# 00footnotetext: Shared first author; Shared last author

### v3

@@ 5. Q26
    • The data was downloaded between January 2023 and May 2023. The urls are collected from Common Crawl archives up to 2023. Common Crawl archives may include URLs from the early days of the internet. Hence, the download / collection timeframe does not match the creation timeframe. Additionally, future users of DCLM-Pool and its subsets will have to download data themselves using our tooling, though the snapshot should not be altered in any way.
@@ 1. Q45
    • Yes. We use HuggingFace datasets for public release.
@@ 2. Q46
    • The dataset will be distributed via HuggingFace.
@@ 1. Q52
    • HuggingFace currently hosts the datasets. The DCLM team will be responsible for maintaining the dataset.
@@ 6. Q57
    • This is the first version of DCLM-Pool and derivative DCLM-baseline dataset. We do not intend to maintain deprecated version of DCLM-Pool. Any deprecation or modification will be announced on our website at https://datacomp.ai/dclm.

### v4

@@ 5. Q26
    • The data was downloaded between January 2023 and May 2023. The urls are collected from Common Crawl archives up to 2023. Common Crawl archives may include URLs from the early days of the internet. Hence, the download / collection timeframe does not necessarily match the creation timeframe. Additionally, future users of DCLM-Pool and its subsets will have to download data themselves using our tooling, though the snapshot should not be altered in any way.
@@ 1. Q45
    • Yes, the datasets are publicly and freely available via both Common Crawl’s S3 bucket and HuggingFace.
@@ 2. Q46
    • The dataset will be distributed via both Common Crawl’s S3 bucket and HuggingFace. Both allow users to make downloads for free.
@@ 1. Q52
    • Common Crawl currently hosts DCLM-Pool and DCLM-baseline, while the competition pools and DCLM-baseline are also available on HuggingFace. The DCLM team will be responsible for maintaining these assets.
@@ 6. Q57
    • This is the first version of DCLM-Pool and derivative DCLM-baseline dataset. We do not intend to maintain deprecated versions of DCLM-Pool. Any deprecation or modification will be announced on our website at https://datacomp.ai/dclm.